OpenStudy (anonymous):

What is the median of the data set? 23 19 13 11 17 19 A. 19 B. 17 C. 12 D. 18

OpenStudy (anonymous):

D 18

OpenStudy (anonymous):

B:

OpenStudy (texaschic101):

first put your numbers in numerical order, and pick the two middle numbers....add them, then divide by 2, and that is your median (middle number)

OpenStudy (anonymous):

The answer is 19

OpenStudy (anonymous):

not 18

OpenStudy (anonymous):

Arranging the given data in ascending order we find 11, 13, 17, 19, 19, 23 here middle terms are 17 and 19 hence nedian = (17+19)/2= 36/2= 18

OpenStudy (anonymous):



OpenStudy (anonymous):

@iknowit how show it..

OpenStudy (texaschic101):

if there is 2 middle numbers, you find the average of the middle numbers by dividing them by 2

OpenStudy (anonymous):

yeah.... so then hang on let me get it ill tell you the answer in a sec.

OpenStudy (anonymous):

@iknowit show that median is 19 and not 18...

OpenStudy (anonymous):

i got 18....

OpenStudy (texaschic101):

correct...it is 18

OpenStudy (anonymous):

Sorry I was wrong :(

OpenStudy (anonymous):

its okay thank you everyone.....hold on

OpenStudy (texaschic101):

no worries...if I had $1 for every mistake I made, I would be rich

OpenStudy (anonymous):

lol...

OpenStudy (anonymous):

thanks though